# Equivalence principles in Weyl transverse gravity

### v1

@@ III.1 Newton equivalence principle
Before going to the more complicated relativistic setting, we first briefly address the weakest formulation of the EP, the Newton EP. It states “In the Newtonian limit, the inertial and gravitational masses of a body are equal” [2]. Since it only deals with the Newtonian limit, it is trivially obeyed in Weyl transverse gravity. We do not see the need to analyse this EP in more detail and refer the interested readers to [2].
III.2 Weak equivalence principles
@@
+~$+\tilde{\nabla}_{\mu}p+\left(\rho+p\right)\tilde{u}^{\nu}\tilde{\nabla}_{\nu}\tilde{u}_{\mu}=\tilde{\nabla}_{\mu}\mathcal{J},$ (17)
@@
The left hand side is proportional to the WTDiff-invariant acceleration of the test particle $\tilde{a}^{\rho}=\tilde{u}^{\nu}\tilde{\nabla}_{\nu}\tilde{u}^{\rho}$, whereas the right hand side is the force acting on the particle. Aside from the force sourced by the gradient of the pressure, there is also a new contribution sourced by the gradient of the energy non-conservation measure $\mathcal{J}$. Therefore, a force-free trajectory of a test particle composed of a perfect fluid in WTDiff-invariant geometry is characterised by the condition $\tilde{h}^{\rho\nu}\tilde{\nabla}_{\nu}\left(\mathcal{J}-p\right)=0$ (the equivalent condition in the Diff-invariant case reads $h^{\rho\nu}\nabla_{\nu}p=0$ since the Diff invariance implies $\mathcal{J}=0$). For such a perfect fluid the divergence of the energy-momentum tensor yields the condition $\tilde{u}^{\nu}\tilde{\nabla}_{\nu}\tilde{u}^{\rho}=0$ and the particle consequently follows a timelike geodesic trajectory in any Weyl frame. Therefore, allowing for non-affine parametrisations, the appropriate WTDiff-invariant geodesic equation reads
@@
For WTDiff-invariant theories, one needs to apply the Geroch-Jang theorem to $\tilde{T}^{\prime}_{\mu\nu}$ (14) whose WTDiff-invariant divergence vanishes as required. Of course, demanding the dominant energy condition for $\tilde{T}^{\prime}_{\mu\nu}$ rather than for $\tilde{T}_{\mu\nu}$ is a stronger requirement. However, equations (13) which are the divergenceless equations for Weyl transverse gravity have $\tilde{T}^{\prime}_{\mu\nu}$ on the right hand side. In other words, it plays the same role as the energy-momentum tensor $T_{\mu\nu}$ in general relativity. Thus, $\tilde{T}^{\prime}_{\mu\nu}$ should be relevant for any application of the energy conditions to WTDiff-invariant gravity, e.g. for the proofs of singularity theorems or for the exclusion of solutions containing closed timelike curves. As an aside, this difference is irrelevant for the null energy conditions, since $\mathcal{J}\tilde{g}_{\mu\nu}\tilde{k}^{\mu}\tilde{k}^{\nu}=0$ for any null vector $\tilde{k}^{\mu}$. With the dominant energy condition satisfied, the Geroch-Jang theorem then ensures the validity of the weak EP for any local, WTDiff-invariant gravitational theory.
@@ III.3 Einstein equivalence principle
Switching from Diff-invariant to WTDiff-invariant gravity does not influence the non-gravitational test physics. In particular, Weyl transformations do not act on the matter fields and the Diff-invariant and WTDiff-invariant laws applying to non-gravitational physics thus have the same form, simply replacing $g_{\mu\nu}$ with $\tilde{g}_{\mu\nu}$ and $\nabla_{\mu}$ with ~$\tilde{\nabla}_{\mu}$ (see the discussion of the geodesic equation in the previous subsection).
@@ IV Gravitational equivalence principles
The weak EP fixes the kinematics of motion in a gravitational field, essentially requiring that the trajectories of the test particles are determined by the connection. The Einstein EP further specifies the behaviour of matter fields in a curved background. However, they are both compatible with essentially any local, Diff- or WTDiff-invariant theory of gravity and do not really constrain the gravitational dynamics in any way. For that, some version of the EP which applies to test gravitational physics is necessary. These EPs present far more stringent constraints and only the Diff-invariant Lanczos-Lovelock theories of gravity are known to be fully compatible with them. In this section, we show that Weyl transverse gravity (as well as the WTDiff-invariant Lanczos-Lovelock theories) also obeys the EP for gravitational test physics.
IV.1 Gravitational weak equivalence principles
@@
A simple criterion for the validity of the gravitational weak EP utilises the Geroch-Jang theorem [5]. However, rather than applying the theorem just to the energy-momentum tensor of the test particle, it also needs to include the perturbation of the gravitational field caused by the presence of the particle (i.e., the effective energy-momentum of its gravitational field). Moreover, one must keep in mind that the geodesic along which the test particle should move lies in the unperturbed spacetime. Splitting the WTDiff-invariant auxiliary metric into the background part $\tilde{g}_{\mu\nu}$ and the perturbation caused by the test particle, $\tilde{\gamma}_{\mu\nu}$, we may similarly split the equations of motion. In the case of Weyl transverse gravity, we obtain the vacuum divergenceless equations for the background metric
@@
Lastly, we must check condition (iii), i.e., that ~$\tilde{\nabla}_{\nu}\tilde{\mathcal{T}}_{\mu}^{\;\>\nu}=0$, where the covariant derivative ~$\tilde{\nabla}_{\nu}$ is defined with respect to the background metric $\tilde{g}_{\mu\nu}$. The gravitational energy-momentum $\tilde{T}^{\text{(g)}}_{\mu\nu}$ is a complicated expression quadratic in the metric perturbation $\tilde{\gamma}_{\mu\nu}$. It is then more convenient to check that ~$\tilde{\nabla}_{\nu}\tilde{\mathcal{G}}_{\mu}^{\;\>\nu}=0$ and use $\tilde{\mathcal{G}}_{\mu\nu}=8\pi G\tilde{\mathcal{T}}_{\mu\nu}$ thanks to equation (21[5].
The divergenceless condition then reads ~$\tilde{\nabla}^{\nu}\left(\tilde{\mathcal{G}}_{\mu\nu}+\Lambda\tilde{\gamma}_{\mu\nu}\right)=0$. Writing it in terms of the metric perturbation $\tilde{\gamma}_{\mu\nu}$ and its derivatives, we obtain
~$\tilde{\nabla}^{\nu}\left(\tilde{\mathcal{G}}_{\mu\nu}+\Lambda\tilde{\gamma}_{\mu\nu}\right)=\frac{1}{2}\tilde{\nabla}^{\nu}\Big{[}2\tilde{\nabla}^{\lambda}\tilde{\nabla}_{(\mu}\tilde{\gamma}_{\nu)\lambda}$
~$-\tilde{\nabla}^{\lambda}\tilde{\nabla}_{\lambda}\tilde{\gamma}_{\mu\nu}+\tilde{g}_{\mu\nu}\left(-\tilde{\nabla}_{\lambda}\tilde{\nabla}_{\rho}\tilde{\gamma}^{\lambda\rho}+\tilde{R}_{\lambda\rho}\tilde{\gamma}^{\lambda\rho}\right)$
@@
where $\tilde{g}_{\mu\nu}$, ~$\tilde{\nabla}_{\mu}$, $\tilde{R}_{\mu\nu}$ denote the corresponding background quantities. We can simplify equation (22) by commuting the derivatives and using the definition of the auxiliary Riemann tensor. Then, we obtain
~$\tilde{\nabla}^{\nu}\left(\tilde{\mathcal{G}}_{\mu\nu}+\Lambda\tilde{\gamma}_{\mu\nu}\right)=\tilde{R}_{\mu}^{\;\>\lambda}\tilde{\nabla}_{\nu}\tilde{\gamma}^{\nu}_{\lambda}+\tilde{\gamma}^{\nu}_{\lambda}\tilde{\nabla}_{\nu}\tilde{R}_{\mu}^{\;\>\lambda}$
@@
In total, $\tilde{\mathcal{T}}_{\mu\nu}$ satisfies all the conditions of the Geroch-Jang theorem. It follows that the test particle moves along a timelike geodesic of the unperturbed auxiliary metric and, consequently, the gravitational weak EP holds. We have shown that, just like general relativity, Weyl transverse gravity incorporates the gravitational weak EP.
Regarding the more general WTDiff-invariant theories, only Lanczos-Lovelock gravity obeys the gravitational weak EP. The proof would be a simple modification of the argument presented for Diff-invariant gravity [5], which has reached the same conclusion. In summary, Weyl transverse gravity and general relativity are the only two metric gravitational theories in four dimensions known to be compatible with the gravitational weak EP (Nordström gravity also satisfies this principle, but it is not metric and, therefore, incompatible with the Einstein EP [5]).
@@ IV.2 Strong equivalence principle
Lastly, the strong EP extends the Einstein EP to include test gravitational physics: “All test fundamental physics (including gravitational physics) is not affected locally by the presence of a gravitational field” [2]. It relates to the Einstein EP in an analogous way as the gravitational weak EP does to the weak EP, extending it to include gravitational fields with a negligible back reaction on the background spacetime. The strong EP can also be phrased as the requirement of local Poincaré invariance of all the test physics, including gravitational physics (e.g. the local behaviour of linearised gravitational waves on a curved background), combined with the validity of the gravitational weak EP [2]. In the previous subsection, we have proven the latter requirement for Weyl transverse gravity. We have also argued in section III that the condition of local Poincaré invariance of matter test physics can be implemented in the WTDiff-invariant setting without any issues (via, e.g. Ehlers criterion). The only remaining issue is the local Poincaré invariance of gravitational test physics, in particular, in regards to weak gravitational waves. However, linearised Weyl transverse gravity is manifestly Poincaré invariant [21, 24, 26], implying local Poincaré invariance of gravitational test physics. The strong EP then does apply to Weyl transverse gravity, as it satisfies all the necessary conditions. Consequently, general relativity and Weyl transverse gravity seem to be the only two known gravitational theories in $4$ spacetime dimensions compatible with the strong EP. The uniqueness of these two theories can be traced to their status as the only two consistent theories of massless, self-interacting gravitons with no additional degrees of freedom [24] (or gauge fixing). Any other propagating gravitational degrees of freedom tend to destroy the gravitational weak EP [5]. For instance, if a non-minimally coupled scalar field is present, the geodesic motion of self-gravitating test particles depends on its local value, breaking the gravitational weak EP.
@@ V Conclusions
We have explored the validity of various formulations of the equivalence principle in WTDiff-invariant theories and, in particular, in Weyl transverse gravity. Our discussion shows how the EPs serve to illuminate the main features of gravity and its coupling to matter. That becomes especially useful in the case of WTDiff-invariant gravity, whose local symmetries and dependence on a nondynamical volume measure lead to behaviour which differs from the usual intuition one develops for Diff-invariant theories.
The weak EP exposes why the geodesic equation needs to be written in term of the Weyl connection $\tilde{\Gamma}^{\mu}_{\;\>\nu\rho}$. Moreover, it points towards the appropriate way for handling the potential local energy-momentum non-conservation in Weyl transverse gravity. The Geroch-Jang theorem which implies the weak EP takes as its input the divergenceless tensor $\tilde{T}^{\prime}_{\mu\nu}$. This tensor also appears as the right hand side of the divergenceless equations for gravitational dynamics. Therefore, it completely determines the motion of matter in a gravitational field. The equations of motion for the matter fields can be further derived from the condition ~$\tilde{\nabla}^{\nu}\tilde{T}^{\prime}_{\mu\nu}=0$. Then, $\tilde{T}^{\prime}_{\mu\nu}$ fulfills all the roles of a locally conserved energy-momentum tensor and we can apparently discard the non-conserved tensor $\tilde{T}_{\mu\nu}$. The potential local energy-momentum non-conservation in WTDiff-invariant gravity then does not appear to have observable consequences (although this conclusion may not apply if, for any reason, ~$\tilde{\nabla}^{\nu}\tilde{T}^{\prime}_{\mu\nu}=0$ does not yield viable equations of motion for the matter fields).
Similarly, the Einstein EP shows that any Diff-invariant theory of matter fields can be rewritten to a WTDiff-invariant one by simply replacing $g_{\mu\nu}$ with $\tilde{g}_{\mu\nu}$ and $\nabla_{\mu}$ with ~$\tilde{\nabla}_{\mu}$. The resulting theory continues to satisfy the Einstein EP, i.e., it remains locally special relativistic.
The EPs for gravitational test physics finally restrict the dynamics. The gravitational weak EP already restricts us to Lanczos-Lovelock theories of gravity (both Diff- and WTDiff- invariant), together with non-metric Nordström gravity [5]. The latter is ruled out by the strong EP. It essentially amounts to requiring that the Einstein EP holds and that the only propagating degrees of freedom of the theory are those associated with a massless graviton. In $4$ spacetime dimensions, the only two theories compatible with the strong EP are then general relativity and Weyl transverse gravity.

### v2

@@ III.1 Newton equivalence principle
Before going to the more complicated relativistic setting, we first briefly address the weakest formulation of the EP, the Newton EP. It states “In the Newtonian limit, the inertial and gravitational masses of a body are equal” [2]. Since it only deals with the Newtonian limit, it is trivially obeyed in WTG. We do not see the need to analyse this EP in more detail and refer the interested readers to [2].
III.2 Weak equivalence principle
@@
$+\tilde{\nabla}_{\mu}p+\left(\rho+p\right)\tilde{u}^{\nu}\tilde{\nabla}_{\nu}\tilde{u}_{\mu}=\tilde{\nabla}_{\mu}\mathcal{J},$ (23)
@@
The left hand side is proportional to the WTDiff-invariant acceleration of the test particle $\tilde{a}^{\rho}=\tilde{u}^{\nu}\tilde{\nabla}_{\nu}\tilde{u}^{\rho}$, whereas the right hand side is the force acting on the particle. Aside from the force sourced by the gradient of the pressure, there is also a new contribution sourced by the gradient of the energy non-conservation measure $\mathcal{J}$. Therefore, a force-free trajectory of a test particle composed of a perfect fluid in WTDiff-invariant geometry is characterised by the condition
$\tilde{h}^{\rho\nu}\tilde{\nabla}_{\nu}\left(\mathcal{J}-p\right)=0.$ (25)
The equivalent condition in the Diff-invariant case reads $h^{\rho\nu}\nabla_{\nu}p=0$ since the Diff invariance implies $\mathcal{J}=0$. Since the divergenceless tensor $\tilde{T}^{\prime}_{\mu\nu}$ for a perfect fluid reads
$\tilde{T}^{\prime}_{\mu\nu}=\left(\rho+\mathcal{J}\right)\tilde{u}_{\mu}\tilde{u}_{\nu}+\left(p-\mathcal{J}\right)\tilde{h}_{\mu\nu},$ (26)
and its effective pressure then equals $p-\mathcal{J}$. Then, the condition for a force-free trajectory requires nothing but the vanishing spatial gradient of the pressure associated with the divergenceless tensor $\tilde{T}^{\prime}_{\mu\nu}$. This tensor is then the relevant measure of the energy-momentum for the purpose of identifying force-free trajectories of test particles.
For a perfect fluid obeying the condition (25) the divergence of the energy-momentum tensor yields the condition $\tilde{u}^{\nu}\tilde{\nabla}_{\nu}\tilde{u}^{\rho}=0$ and the particle consequently follows a timelike geodesic trajectory in any Weyl frame. Therefore, allowing for non-affine parametrisations, the appropriate WTDiff-invariant geodesic equation reads
@@
For WTDiff-invariant theories, one needs to apply the Geroch-Jang theorem to $\tilde{T}^{\prime}_{\mu\nu}$ (14) whose WTDiff-invariant divergence vanishes as required. Of course, demanding the dominant energy condition for $\tilde{T}^{\prime}_{\mu\nu}$ rather than for $\tilde{T}_{\mu\nu}$ is a stronger requirement. However, equations (13) which are the divergenceless equations for WTG have $\tilde{T}^{\prime}_{\mu\nu}$ on the right hand side. In other words, it plays the same role as the energy-momentum tensor $T_{\mu\nu}$ in GR. Thus, $\tilde{T}^{\prime}_{\mu\nu}$ should be relevant for any application of the energy conditions to WTDiff-invariant gravity, e.g. for the proofs of singularity theorems or for the exclusion of solutions containing closed timelike curves. As an aside, this difference is irrelevant for the null energy conditions, since $\mathcal{J}\tilde{g}_{\mu\nu}\tilde{k}^{\mu}\tilde{k}^{\nu}=0$ for any null vector $\tilde{k}^{\mu}$. With the dominant energy condition satisfied, the Geroch-Jang theorem then ensures the validity of the weak EP for any local, WTDiff-invariant gravitational theory.
@@ III.3 Einstein equivalence principle
Switching from Diff-invariant to WTDiff-invariant gravity does not influence the non-gravitational test physics. In particular, Weyl transformations do not act on the matter fields and the Diff-invariant and WTDiff-invariant laws applying to non-gravitational physics thus have the same form, simply replacing $g_{\mu\nu}$ with $\tilde{g}_{\mu\nu}$ and $\nabla_{\mu}$ with $\tilde{\nabla}_{\mu}$ (see the discussion of the geodesic equation in the previous subsection).
@@
III.4 Equivalence principles and local energy-momentum non-conservation
Passing the tests for the EPs for non-gravitational test physics serves primarily as a sanity check for WTDiff-invariant gravity. Beyond that, it also illuminates the issue of the possible local energy-momentum non-conservation proposals in these theories. While we touched on this question in the previous subsections, we can now provide a synthesis of the insights from the different EPs.
To begin, we stress again that the equations of motion of WTDiff-invariant gravities depend only on the traceless part of the energy-momentum tensor. Even applying the Bianchi identities to obtain divergenceless equations, the right hand side is given by the divergenceless tensor $\tilde{T}^{\prime}_{\mu\nu}$ defined by equation (14). Then, the possible local energy-momentum non-conservation cannot be actually noticed at the level of the gravitational equations of motion. Then, the kinematics of matter fields in a WTDiff-invariant spacetime, constrained by the EPs, is a natural place to look for signatures of the local energy-momentum non-conservation.
In the context of the weak EP, we have shown that the relevant tensor for the Geroch-Jang theorem is again the divergenceless $\tilde{T}^{\prime}_{\mu\nu}$. Likewise, the simpler derivation from the perfect fluid tensor imposes the condition of the vanishing pressure gradient on the effective pressure term $p-\mathcal{J}$, associated with the divergenceless tensor $\tilde{T}^{\prime}_{\mu\nu}$, which for a perfect fluid follows equation (26).
In conclusion, the weak EP only cares about $\tilde{T}^{\prime}_{\mu\nu}$. This outcome is consistent with the fact that the weak EP can be heuristically derived from the gravitational equations of motion, by considering a point-like source for them. Since the equations of motion only depend on $\tilde{T}^{\prime}_{\mu\nu}$, it follows that the same applies to the weak EP.
The Einstein EP does not only concern the geodesic equation but also the equations of motion of the matter fields. Since the local energy-momentum (non-)conservation is determined by the matter equations of motion (assuming one can independently identify the energy-momentum tensor [^3] This issue is in general very subtle, since theories with energy-momentum non-conservation generically lack a Lagrangian description [42, 43, 23].), the Einstein EP is sensitive to it. Nevertheless, local energy-momentum non-conservation does not signal any breaking of the Einstein EP. One only requires that the non-conservation (on the level of fundamental physics as required by the Einstein EP) does not locally depend on the spacetime curvature or on the background volume measure $\omega$. This condition only amounts to minimal coupling of the matter fields to gravity, which is a general requirement for the Einstein EP to hold. Assuming minimal coupling, the Ehlers criterion goes through by matching the solutions of the flat spacetime matter equations of motion with the curved spacetime solutions in a suitable frame. In this sense, the local energy-momentum (non-)conservation introduces no to the Einstein EP.
For instance, a perfect fluid model with energy-momentum non-conservation we applied to discuss the weak equivalence principle passes the Ehlers criterion, provided that $\mathcal{J}$ does not depend on the Riemann tensor or on $\omega$. It is easy to see that the flat spacetime and curved spacetime solutions can be matched in this case. A special case of such a perfect fluid is the continuous spontaneous localisation model [44, 45] suggested in the cosmological context as a possible source for the acceleration of the universe [42]. This proposal involves no violation of the Einstein EP on the level of classical physics [^4] A later proposal of energy dispersion due to granular structure of the spacetime [43] represents a special case, as it involves point-like spinning particles. These are known to generically violate even the weak EP [2]. The EP violation then does not depend on the local energy non-conservation occurring in this model..
@@ IV Gravitational equivalence principles
The weak EP fixes the kinematics of motion in a gravitational field, essentially requiring that the trajectories of the test particles are determined by the connection. The Einstein EP further specifies the behaviour of matter fields in a curved background. However, they are both compatible with essentially any local, Diff- or WTDiff-invariant theory of gravity and do not really constrain the gravitational dynamics in any way. For that, some version of the EP which applies to test gravitational physics is necessary. These EPs present far more stringent constraints and only the Diff-invariant Lanczos-Lovelock theories of gravity are known to be compatible even with the gravitational weak EP. In this section, we show that WTG (as well as the WTDiff-invariant Lanczos-Lovelock theories) also obeys the gravitational weak EP and that WTG incorporates even the more stringent strong EP.
IV.1 Gravitational weak equivalence principle
@@
A simple criterion for the validity of the gravitational weak EP utilises the Geroch-Jang theorem [5]. However, rather than applying the theorem just to the energy-momentum tensor of the test particle, it also needs to include the perturbation of the gravitational field caused by the presence of the particle (i.e., the effective energy-momentum of its gravitational field). Moreover, one must keep in mind that the geodesic along which the test particle should move lies in the unperturbed spacetime. Splitting the WTDiff-invariant auxiliary metric into the background part $\tilde{g}_{\mu\nu}$ and the perturbation caused by the test particle, $\tilde{\gamma}_{\mu\nu}$, we may similarly split the equations of motion. In the case of WTG, we obtain the vacuum divergenceless equations for the background metric
@@
Lastly, we must check condition (iii), i.e., that $\tilde{\nabla}_{\nu}\tilde{\mathcal{T}}_{\mu}^{\;\>\nu}=0$, where the covariant derivative $\tilde{\nabla}_{\nu}$ is defined with respect to the background metric $\tilde{g}_{\mu\nu}$. The gravitational energy-momentum $\tilde{T}^{\text{(g)}}_{\mu\nu}$ is a complicated expression quadratic in the metric perturbation $\tilde{\gamma}_{\mu\nu}$. It is then more convenient to check that $\tilde{\nabla}_{\nu}\tilde{\mathcal{G}}_{\mu}^{\;\>\nu}=0$ and use $\tilde{\mathcal{G}}_{\mu\nu}=8\pi G\tilde{\mathcal{T}}_{\mu\nu}$ thanks to equation (29[5].
The divergenceless condition then reads $\tilde{\nabla}^{\nu}\left(\tilde{\mathcal{G}}_{\mu\nu}+\Lambda\tilde{\gamma}_{\mu\nu}\right)=0$. Writing it in terms of the metric perturbation $\tilde{\gamma}_{\mu\nu}$ and its derivatives, we obtain
$\tilde{\nabla}^{\nu}\left(\tilde{\mathcal{G}}_{\mu\nu}+\Lambda\tilde{\gamma}_{\mu\nu}\right)=\frac{1}{2}\tilde{\nabla}^{\nu}\Big[2\tilde{\nabla}^{\lambda}\tilde{\nabla}_{(\mu}\tilde{\gamma}_{\nu)\lambda}$
$-\tilde{\nabla}^{\lambda}\tilde{\nabla}_{\lambda}\tilde{\gamma}_{\mu\nu}+\tilde{g}_{\mu\nu}\left(-\tilde{\nabla}_{\lambda}\tilde{\nabla}_{\rho}\tilde{\gamma}^{\lambda\rho}+\tilde{R}_{\lambda\rho}\tilde{\gamma}^{\lambda\rho}\right)$
@@
where $\tilde{g}_{\mu\nu}$, $\tilde{\nabla}_{\mu}$, $\tilde{R}_{\mu\nu}$ denote the corresponding background quantities. We can simplify equation (30) by commuting the derivatives and using the definition of the auxiliary Riemann tensor. Then, we obtain
$\tilde{\nabla}^{\nu}\left(\tilde{\mathcal{G}}_{\mu\nu}+\Lambda\tilde{\gamma}_{\mu\nu}\right)=\tilde{R}_{\mu}^{\;\>\lambda}\tilde{\nabla}_{\nu}\tilde{\gamma}^{\nu}_{\lambda}+\tilde{\gamma}^{\nu}_{\lambda}\tilde{\nabla}_{\nu}\tilde{R}_{\mu}^{\;\>\lambda}$
@@
In total, $\tilde{\mathcal{T}}_{\mu\nu}$ satisfies all the conditions of the Geroch-Jang theorem. It follows that the test particle moves along a timelike geodesic of the unperturbed auxiliary metric and, consequently, the gravitational weak EP holds. We have shown that, just like GR, WTG incorporates the gravitational weak EP.
Regarding the more general WTDiff-invariant theories, only Lanczos-Lovelock gravity obeys the gravitational weak EP. The proof would be a simple modification of the argument presented for Diff-invariant gravity [5], which has reached the same conclusion. In summary, WTG and GR are the only two metric gravitational theories in four dimensions known to be compatible with the gravitational weak EP (Nordström gravity also satisfies this principle, but it is not metric and, therefore, incompatible with the Einstein EP [5]).
@@ IV.2 Strong equivalence principle
Lastly, the strong EP extends the Einstein EP to include test gravitational physics: “All test fundamental physics (including gravitational physics) is not affected locally by the presence of a gravitational field” [2]. It relates to the Einstein EP in an analogous way as the gravitational weak EP does to the weak EP, extending it to include gravitational fields with a negligible back reaction on the background spacetime. The strong EP can also be phrased as the requirement of local Poincaré invariance of all the test physics, including gravitational physics (e.g. the local behaviour of linearised gravitational waves on a curved background), combined with the validity of the gravitational weak EP [2]. In the previous subsection, we have proven the latter requirement for WTG. We have also argued in section III that the condition of local Poincaré invariance of matter test physics can be implemented in the WTDiff-invariant setting without any issues (via, e.g. Ehlers criterion). The only remaining issue is the local Poincaré invariance of gravitational test physics, in particular, in regards to weak gravitational waves. However, linearised WTG is manifestly Poincaré invariant [18, 21, 23], implying local Poincaré invariance of gravitational test physics. The strong EP then does apply to WTG, as it satisfies all the necessary conditions. Consequently, GR and WTG seem to be the only two known gravitational theories in $4$ spacetime dimensions compatible with the strong EP. The uniqueness of these two theories can be traced to their status as the only two consistent theories of massless, self-interacting gravitons with no additional degrees of freedom [21] (or gauge fixing). Any other propagating gravitational degrees of freedom tend to destroy the gravitational weak EP [5]. For instance, if a non-minimally coupled scalar field is present, the geodesic motion of self-gravitating test particles depends on its local value, breaking the gravitational weak EP.
@@ V Conclusions
We have explored the validity of various formulations of the equivalence principle in WTDiff-invariant theories and, in particular, in WTG. Our discussion shows how the EPs serve to illuminate the main features of gravity and its coupling to matter. That becomes especially useful in the case of WTDiff-invariant gravity, whose local symmetries and dependence on a nondynamical volume measure lead to behaviour which differs from the usual intuition one develops for Diff-invariant theories.
The weak EP exposes why the geodesic equation needs to be written in term of the Weyl connection $\tilde{\Gamma}^{\mu}_{\;\>\nu\rho}$. Moreover, it points towards the appropriate way for handling the potential local energy-momentum non-conservation in WTG. The Geroch-Jang theorem which implies the weak EP takes as its input the divergenceless tensor $\tilde{T}^{\prime}_{\mu\nu}$. This tensor also appears as the right hand side of the divergenceless equations for gravitational dynamics. Therefore, it completely determines the motion of matter in a gravitational field. The equations of motion for the matter fields can be further derived from the condition $\tilde{\nabla}^{\nu}\tilde{T}^{\prime}_{\mu\nu}=0$. Then, on the level of the gravitational dynamics, $\tilde{T}^{\prime}_{\mu\nu}$ fulfills all the roles of a locally conserved energy-momentum tensor. Nevertheless, the non-conserved energy-momentum tensor $\tilde{T}_{\mu\nu}$ can still be significant if it corresponds to the definitions of energy, momenta and stresses that are “natural” from the perspective of the matter field dynamics (see, e.g. [44, 45] for an example of such a matter theory).
Similarly, the Einstein EP shows that any Diff-invariant theory of matter fields can be rewritten to a WTDiff-invariant one by simply replacing $g_{\mu\nu}$ with $\tilde{g}_{\mu\nu}$ and $\nabla_{\mu}$ with $\tilde{\nabla}_{\mu}$. The resulting theory continues to satisfy the Einstein EP, i.e., it remains locally special relativistic.
The EPs for gravitational test physics finally restrict the dynamics. The gravitational weak EP already restricts us to Lanczos-Lovelock theories of gravity (both Diff- and WTDiff- invariant), together with non-metric Nordström gravity [5]. The former has been shown to fail the strong EP [48]. The latter is also trivially ruled out by the strong EP. It essentially amounts to requiring that the Einstein EP holds and that the only propagating degrees of freedom of the theory are those associated with a massless graviton. In $4$ spacetime dimensions, the only two theories compatible with the strong EP are then GR and WTG.